Mr. Wilson invested money in two accounts. His total investment was $30,000. If one account pays 6% in interest and the other pays 8% in interest, how much does he have in each account if he earned a total of $1,920 in interest in 1 year?
He invested $
in the 6% account and $
in the 8% account.
I tried this question twice and received two different answers, just needing a formula

Answer by mananth(16946)

Mr. Wilson invested money in two accounts. His total investment was $30,000. If one account pays 6% in interest and the other pays 8% in interest, how much does he have in each account if he earned a total of $1,920 in interest in 1 year?
He invested $
in the 6% account and $
in the 8% account.

Part I 6.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y =
30000
Interest----- 1920

Part I 6.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 30000 -------------1
Interest on both investments
6.00% x + 8.00% y= 1920
Multiply by 100
6 x + 8 y= 192000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -180000.00
Add this to (2)
0 x 2 y= 12000
divide by 2
y = 6000
Part I 6.00% $ 24000
Part II 8.00% $ 6000

CHECK
24000 --------- 6.00% ------- 1440.00
6000 ------------- 8.00% ------- 480.00
Total -------------------- 1920.00

m.ananth@hotmail.ca